OpenStudy (anonymous):

What is the y-intercept for the equation 5x-2y=-4?

OpenStudy (anonymous):

The y-intercept is what y is when x = 0. So plug in x for 0 in the equation and solve for y.

OpenStudy (anonymous):

5x-2y=-4 5(0) - 2y = -4 0 - 2y = -4 -2y = -4 Solve it from here.

OpenStudy (anonymous):

-2y = -4 Divide -2 to both sides: y = 2

OpenStudy (haseeb96):

for y intercept put x =0 in the given equation